# Quasi-likelihood inference for SDE with mixed-effects observed at high frequency

### v1

@@ A.2.1. Fixed-effect parameter
$|\pi_{i,j-1}|\leq C(1+|Y_{i}(t_{j-1})|^{C}).$ (A.35)
Further, we write $E^{i,j-1}[\cdot]$ for the conditional expectation $E[\cdot|\mathcal{F}_{i,t_{j-1}}]$.
@@
for sufficiently integrable $(\mathcal{F}_{i,t})_{t\leq T}$-adapted process $(\alpha_{i},\beta_{i})$ and also applying Itô’s formula, the second term in the right-hand side of (A.36) equals
@@
This shows that $\overline{\Delta}_{11}^{r,3,1}=O_{p}(\sqrt{N/n})=o_{p}(1)$.
The term $\overline{\Delta}_{11}^{r,3,2}$ can be handled similarly: as before, through the compensation of $c_{i,j-1}^{-1}\mu_{ij}^{\star}z_{ij}$, we obtain
$\overline{\Delta}_{11}^{r,3,2}$ $\leq O_{p}\left(\sqrt{\frac{N}{n}}\right)+\frac{1}{N}\sum_{i=1}^{N}|\overline{\mathfrak{m}}_{i}|\left|\frac{1}{n}\sum_{j}c_{i,j-1}^{-1}E^{i,j-1}[\mu_{ij}^{\star}z_{ij}]\right|\sqrt{N}$ (A.39)
Plugging-in the expression (A.10) of $\mu_{ij}^{\star}$ and then handling it as in the case of (A.36), we obtain $\overline{\Delta}_{11}^{r,3,2}=O_{p}(\sqrt{N/n})=o_{p}(1)$.
Thus we have obtained $\Delta_{11}=\Delta_{11}^{\star}+o_{p}(1)$. Let
@@
$\Delta_{11}$ $=\frac{1}{2\sqrt{N}}\sum_{i=1}^{N}\Bigg{\{}\frac{1}{\sqrt{n}}\sum_{j}g_{i,j-1}(z_{ij}^{2}-1)$
@@
Denote by $\sum_{i=1}^{N}\zeta_{i}$ the first on the rightmost side.
We will show that
@@
$=-\frac{1}{2\sqrt{nN}}\sum_{j}E\left[E\big{[}(\overline{g}_{i}-E[\overline{g}_{i}])(z_{ij}^{2}-1)\middle|(\tau_{i},\varphi_{i})\big{]}\right]$ (A.44)
Conditional on $\xi_{i}=(\tau_{i},\varphi_{i})$, the variables $\overline{g}_{i}$ and $z_{ij}$ are functionals of $\{Y_{i}(0),w_{i}\}$; again, recall that $\xi_{i}$ and $\{Y_{i}(0),w_{i}\}$ are independent. Itô’s formula gives the expression
@@
By applying the representation theorem [23, Theorem 1.1.3] conditional on $\xi_{i}$, there exists an $(\mathcal{F}_{i,t})$-adapted $L^{2}$-process $\psi_{i}=\psi_{i,T}$ such that
@@
By the martingale representation theorem [19, Theorem III.4.33 a)], there exists an $(\mathcal{F}_{i,t})_{t\leq T}$-predictable matrix-valued process $\psi^{\prime}_{i}$ for which the right-hand side of (A.48) is written as
@@
$|E[\zeta_{i}]|$ $=\left|-\frac{1}{2\sqrt{nN}}\sum_{j}\frac{2}{h}E\left[\int_{j}\int_{t_{j-1}}^{s}\psi^{\prime}_{i}(u)duds\right]\right|\lesssim\frac{1}{\sqrt{nN}}.$ (A.50)
Therefore,
$\Delta_{11}=\sum_{i=1}^{N}(\zeta_{i}-E[\zeta_{i}])+o_{p}(1)=:\sum_{i=1}^{N}\widetilde{\zeta}_{i}+o_{p}(1).$ (A.51)
It is easy to see that for $K>2$, $\sum_{i=1}^{N}E[|\widetilde{\zeta}_{i}|^{K}]\lesssim N^{1-K/2}\to 0$. The convergence of the covariance $Q_{11}:=\sum_{i=1}^{N}E[\widetilde{\zeta}_{i}^{\otimes 2}]$ remains to be studied.
Observe that $Q_{11}=\sum_{i=1}^{N}E[\zeta_{i}^{\otimes 2}]+O(n^{-1})$ by (A.43). Since $(\tau_{i},\varphi_{i},X_{i},Y_{i})$, $i\geq 1$, are i.i.d., we have
$\sum_{i=1}^{N}E[\zeta_{i}^{\otimes 2}]$ $=\frac{1}{4}E\left[\{M_{11,1}-\overline{g}_{i}M_{11,0}\}^{\otimes 2}\right],$ (A.52)
where $M_{11,1}:=n^{-1/2}\sum_{j}g_{i,j-1}(z_{ij}^{2}-1)$ and $M_{11,0}:=n^{-1/2}\sum_{j}(z_{ij}^{2}-1)$. We can apply [19, Theorem IX.7.28] for the step process
@@
$\sim\sqrt{2}\begin{pmatrix}T^{-1}\int_{0}^{T}g(X_{1}(t))^{\otimes 2}dt&\text{sym.}\\ \overline{g}_{1}&1\end{pmatrix}^{1/2}Z$ (A.54)
defined on a suitable extended probability space, where $Z\sim N_{p_{\eta_{1}}}(0,I_{p_{\eta_{1}}})$ is independent of $\mathcal{F}$; to apply the theorem, we set $\overline{w}_{1}$ to be a reference continuous local martingale, and make use of the representation theorem [19, Theorem III.4.33 a)] as before to verify the asymptotic orthogonality condition [19, IX.7.31]. Since $\overline{g}_{i}$ is $\mathcal{F}$-measurable, the stability guarantees the joint weak convergence
@@
is obvious so that in particular, the sequence $\{(\overline{M}_{11}(T),\overline{g}_{1})\}_{n}$ is uniformly integrable. By (A.52),
@@
$=\frac{1}{2}E\left[\frac{1}{T}\int_{0}^{T}g(X_{1}(t))^{\otimes 2}dt-\overline{g}_{1}^{\otimes 2}\right]=:Q_{11,0},$ (A.58)
where the limit is positive definite under Assumption 2.2(2). We have arrived at
$\Delta_{11}\xrightarrow{\mathcal{L}}N_{p_{\eta}}(0,Q_{11,0}).$ (A.59)
@@
$\Gamma_{11}$ $=\frac{1}{2N}\sum_{i=1}^{N}\Bigg{\{}\frac{1}{n}\sum_{j}(\partial_{\eta}^{2}\log S)_{i,j-1}+Q_{i}^{-1}\frac{1}{n}\sum_{j}(\partial_{\eta}^{2}(S^{-1}))_{i,j-1}y_{ij}^{2}$
${}\qquad-Q_{i}^{-2}\bigg{(}\frac{1}{n}\sum_{j}(\partial_{\eta}(S^{-1}))_{i,j-1}y_{ij}^{2}\bigg{)}^{\otimes 2}\Bigg{\}}$
@@
Finally, we note that, based on what we have seen so far it is easy to show that
@@ A.2.2. Random-effect distribution
Recall the definition $\mathbb{H}_{12}(\vartheta_{1})$ in (3.8). With a slight abuse of notation, we let $\mathbb{H}_{12}(\theta_{\tau}):=\mathbb{H}_{12}(\widehat{\eta},\theta_{\tau})$ and write
@@
$\widehat{\tau}_{i}$ $=\frac{1}{n}\sum_{j}\left(S^{-1}_{i,j-1}+\frac{1}{\sqrt{nN}}\int_{0}^{1}\partial_{\eta}(S^{-1})_{i,j-1}\left(\eta_{0})+s(\widehat{\eta}-\eta_{0})\right)ds\,[\widehat{u}_{\eta}]\right)\,y_{ij}^{2}$
$=\frac{1}{n}\sum_{j}\left(S^{-1}_{i,j-1}+\frac{1}{\sqrt{nN}}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\right)\,y_{ij}^{2}$
$=\frac{1}{n}\sum_{j}(\tau_{i}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij})+\frac{1}{\sqrt{nN}}\frac{1}{n}\sum_{j}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\,y_{ij}^{2}$
$=\tau_{i}+\sqrt{h}\,\overline{\mathfrak{m}}_{i}+\frac{1}{\sqrt{nN}}\frac{1}{n}\sum_{j}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\,y_{ij}^{2}.$ (A.72)
Hence, 222 Recalling (A.28), we also have for any $\mathsf{c}>0$, $\max_{i\leq N}|h^{\mathsf{c}-1/2}(\widehat{\tau}_{i}-\tau_{i})|\lesssim(1+N^{-1/2}|\widehat{\eta}|)\max_{i\leq N}|h^{\mathsf{c}}\,\overline{\mathfrak{m}}_{i}|=O_{p}(1)o_{p}(1)=o_{p}(1)$.
$\left|\sqrt{n}(\widehat{\tau}_{i}-\tau_{i})\right|\lesssim|\overline{\mathfrak{m}}_{i}|+\frac{|\widehat{u}_{\eta}|}{\sqrt{N}}\frac{1}{n}\sum_{j}y_{ij}^{2}\lesssim|\overline{\mathfrak{m}}_{i}|\left(1+\frac{|\widehat{u}_{\eta}|}{\sqrt{N}}\right).$ (A.73)
@@
Take $K>0$ large enough to ensure that $\sup_{n,N}P[A_{12,1}^{c}]<\epsilon^{\prime}$. With this $K$, by (A.73) and the Markov inequality we have for any $L\geq 2$,
@@
Finally, by Assumption 2.4, it is easy to see that the last term in the right-hand side of (A.2.2) can be bounded by
$P\left[\sqrt{\frac{N}{n}}\,\frac{1}{N}\sum_{i=1}^{N}|\overline{\mathfrak{m}}_{i}|>\epsilon\right]\lesssim\sqrt{\frac{N}{n}}\to 0.$ (A.77)
@@ A.2.3. Joint asymptotic normality
Having obtained the asymptotic linear representations (A.62) and (A.69), it remains to show the asymptotic orthogonality, hence the asymptotic independence of $\widehat{\eta}$ and $\widehat{\theta}_{\tau}$: we only need to observe that the covariance tends to 0. By what we have seen in Sections A.2.1 and A.2.2, in particular the stochastic expansions (A.63) and (A.69), the computation of the covariance of the $\Delta_{11}$ and $\Delta_{12}$ boils down to that of
${\rm cov}[\zeta_{i},\zeta_{i}^{\prime}]$ $=\frac{1}{2}\,{\rm cov}\left[M_{11,1}-\overline{g}_{1}M_{11,0},\,\partial_{\theta_{\tau}}\log f(\tau_{1};\theta_{\tau,0})\right]$ (A.78)
As in the $\mathcal{F}$-stable convergence argument of $(M_{11,0},M_{11,1})$ combined with the uniform integrability we have used to obtain the limit of the expectation in (A.52), it can be seen that $\lim_{n,N}{\rm cov}[\zeta_{i},\zeta_{i}^{\prime}]=0$ ($p_{\eta}\times p_{\tau}$-zero matrix). This completes the proof of Theorem 3.2.
@@
To study the asymptotic behavior of $\mathbb{H}_{2}(\vartheta_{2})$ of (3.17), we will proceed as follows.
  1. (1)
@@
    for specific leading terms $M_{i,T}$ and $v_{i,T}$ and some remainder (negligible) terms $\mathsf{R}^{M}_{i}$ and $\mathsf{R}^{v}_{i}$.
  2. (2)
@@
    $\mathbb{H}^{\star}_{2}(\vartheta_{2})$ $:=\sum_{i}\log\phi_{p_{\varphi}}\left(M_{i,T}^{-1}v_{i,T};\,\mu,\,M_{i,T}^{-1}+\Sigma\right)$ (A.80)

### v2

@@ A.2.1. Fixed-effect parameter
$|\pi_{i,j-1}|\leq C(1+|Y_{i,j-1}|^{C}).$ (A.37)
Further, we will write $E^{i,j-1}[\cdot]$ for the conditional expectation $E[\cdot|\mathcal{F}_{i,t_{j-1}}]$.
@@
for sufficiently integrable $(\mathcal{F}_{i,t})_{t\leq T}$-adapted process $(\alpha_{i},\beta_{i})$ and also applying Itô’s formula, we can deduce that the second term in the right-hand side of (A.38) equals
@@
We conclude that $\overline{\Delta}_{11}^{r,3,1}=O_{p}(\sqrt{N/n})=o_{p}(1)$.
The term $\overline{\Delta}_{11}^{r,3,2}$ can be handled similarly: through the compensation of $c_{i,j-1}^{-1}\mu_{ij}^{\star}z_{ij}$ as before, we obtain
$\overline{\Delta}_{11}^{r,3,2}$ $\leq O_{p}\left(\sqrt{\frac{N}{n}}\right)+\frac{1}{N}\sum_{i=1}^{N}|\overline{\mathfrak{m}}_{i}|\left|\frac{1}{n}\sum_{j}c_{i,j-1}^{-1}E^{i,j-1}[\mu_{ij}^{\star}z_{ij}]\right|\sqrt{N}.$ (A.41)
Plugging-in the expression (A.13) of $\mu_{ij}^{\star}$ and then handling it as in the case of (A.38), we get $\overline{\Delta}_{11}^{r,3,2}=O_{p}(\sqrt{N/n})=o_{p}(1)$.
Thus we have obtained $\Delta_{11}=\Delta_{11}^{\star}+o_{p}(1)$, hence it remains to study $\Delta_{11}^{\star}$. Let
@@
$\Delta_{11}^{\star}$ $=\frac{1}{2\sqrt{N}}\sum_{i=1}^{N}\Bigg\{\frac{1}{\sqrt{n}}\sum_{j}g_{i,j-1}(z_{ij}^{2}-1)$
@@
Denote by $\sum_{i=1}^{N}\zeta_{i}$ the first term on the rightmost side.
Toward the central limit theorem for $\Delta_{11}^{\star}$, we will show that
@@
$=-\frac{1}{2\sqrt{nN}}\sum_{j}E\left[E\big[(\overline{g}_{i}-E[\overline{g}_{i}])(z_{ij}^{2}-1)\middle|(\tau_{i},\varphi_{i})\big]\right].$ (A.46)
Conditional on $(\tau_{i},\varphi_{i})$, the variables $\overline{g}_{i}$ and $z_{ij}$ are functionals of $\{Y_{i}(0),w_{i}\}$; recall that $(\tau_{i},\varphi_{i})$ and $\{Y_{i}(0),w_{i}\}$ are independent. Itô’s formula gives the expression
@@
By applying the representation theorem [Nua06, Theorem 1.1.3] conditional on $(\tau_{i},\varphi_{i})$, there exists an $(\mathcal{F}_{i,t})$-adapted $L^{2}$-process $\psi_{i}=\psi_{i,T}$ such that
@@
By the martingale representation theorem [JacShi03, Theorem III.4.33 a)], there exists an $(\mathcal{F}_{i,t})_{t\leq T}$-predictable matrix-valued process $\psi^{\prime}_{i}$ for which the right-hand side of (A.50) is written as
@@
$|E[\zeta_{i}]|$ $=\left|-\frac{1}{2\sqrt{nN}}\sum_{j}\frac{2}{h}E\left[\int_{j}\int_{t_{j-1}}^{s}\psi^{\prime}_{i}(u)duds\right]\right|\lesssim\frac{1}{\sqrt{nN}}\sum_{j=1}^{n}\frac{1}{h}h^{2}=\frac{1}{\sqrt{nN}}.$ (A.52)
Now (A.45) leads to
$\Delta_{11}^{\star}=\sum_{i=1}^{N}(\zeta_{i}-E[\zeta_{i}])+o_{p}(1)=:\sum_{i=1}^{N}\widetilde{\zeta}_{i}+o_{p}(1).$ (A.53)
It is easy to see that for $K>2$, $\sum_{i=1}^{N}E[|\widetilde{\zeta}_{i}|^{K}]\lesssim N^{1-K/2}\to 0$, so that the Lyapunov condition holds. The convergence of the covariance $Q_{11}:=\sum_{i=1}^{N}E[\widetilde{\zeta}_{i}^{\otimes 2}]$ remains to be studied.
Observe that $Q_{11}=\sum_{i=1}^{N}E[\zeta_{i}^{\otimes 2}]+O(n^{-1})$ by (A.45). Since $(\tau_{i},\varphi_{i},Y_{i})$, $i\geq 1$, are i.i.d., we have
$\sum_{i=1}^{N}E[\zeta_{i}^{\otimes 2}]$ $=\frac{1}{4}E\left[\{M_{11,1}-\overline{g}_{1}M_{11,0}\}^{\otimes 2}\right],$ (A.54)
where $M_{11,1}:=n^{-1/2}\sum_{j}g_{i,j-1}(z_{ij}^{2}-1)$ and $M_{11,0}:=n^{-1/2}\sum_{j}(z_{ij}^{2}-1)$. We can apply [JacShi03, Theorem IX.7.28] for the step process
@@
$\sim\sqrt{2}\begin{pmatrix}T^{-1}\int_{0}^{T}g_{1}(t)^{\otimes 2}dt&\text{sym.}\\ \overline{g}_{1}&1\end{pmatrix}^{1/2}\mathsf{Z}$ (A.56)
defined on a suitable extended probability space, where $\mathsf{Z}\sim N_{p_{\eta_{1}}}(0,I_{p_{\eta_{1}}})$ is independent of $\mathcal{F}$; to apply the theorem, we set $\overline{w}_{1}$ to be a reference continuous local martingale, and make use of the representation theorem [JacShi03, Theorem III.4.33 a)] as before to verify the asymptotic orthogonality condition [JacShi03, IX.7.31]. Since $\overline{g}_{i}$ is $\mathcal{F}$-measurable, the $\mathcal{F}$-stability of the convergence $\overline{M}_{11}(\cdot)\xrightarrow{\mathcal{L}}\overline{M}_{11,\infty}(\cdot)$ guarantees the joint weak convergence
@@
for any $K>0$ is obvious. Hence, the sequence $(\{M_{11,1}-\overline{g}_{1}M_{11,0}\}^{\otimes 2})_{n}$ is uniformly integrable, which ensures the convergence of moments. By (A.54),
@@
$=\frac{1}{2}E\left[\frac{1}{T}\int_{0}^{T}g_{1}(t)^{\otimes 2}dt-\overline{g}_{1}^{\otimes 2}\right]=:Q_{11,0},$ (A.60)
where the limit is positive definite under Assumption 2.2(2). Thus, we have arrived at
$\Delta_{11}=\Delta_{11}^{\star}+o_{p}(1)\xrightarrow{\mathcal{L}}N_{p_{\eta}}(0,Q_{11,0}).$ (A.61)
@@
$\Gamma_{11}$ $=\frac{1}{2N}\sum_{i=1}^{N}\Bigg\{\frac{1}{n}\sum_{j}(\partial_{\eta}^{2}\log S)_{i,j-1}+Q_{i}^{-1}\frac{1}{n}\sum_{j}(\partial_{\eta}^{2}(S^{-1}))_{i,j-1}\overline{y}_{ij}^{2}$
${}\qquad-Q_{i}^{-2}\bigg(\frac{1}{n}\sum_{j}(\partial_{\eta}(S^{-1}))_{i,j-1}\overline{y}_{ij}^{2}\bigg)^{\otimes 2}\Bigg\}$
@@
Finally, we note that, based on what we have seen so far, it is easy to show that
@@ A.2.2. Random-effect distribution
Recall the definition $\mathbb{H}_{12}(\vartheta_{1})$ in (3.8). Let $\mathbb{H}_{12}(\theta_{\tau}):=\mathbb{H}_{12}(\widehat{\eta},\theta_{\tau})$ with a slight abuse of notation, and write
@@
$\widehat{\tau}_{i}$ $=\frac{1}{n}\sum_{j}\left(S^{-1}_{i,j-1}+\frac{1}{\sqrt{nN}}\int_{0}^{1}\partial_{\eta}(S^{-1})_{i,j-1}\left(\eta_{0})+s(\widehat{\eta}-\eta_{0})\right)ds\,[\widehat{u}_{\eta}]\right)\,\overline{y}_{ij}^{2}$
$=\frac{1}{n}\sum_{j}\left(S^{-1}_{i,j-1}+\frac{1}{\sqrt{nN}}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\right)\,\overline{y}_{ij}^{2}$
$=\frac{1}{n}\sum_{j}(\tau_{i}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij})+\frac{1}{\sqrt{nN}}\frac{1}{n}\sum_{j}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\,\overline{y}_{ij}^{2}$
$=\tau_{i}+\sqrt{h}\,\overline{\mathfrak{m}}_{i}+\frac{1}{\sqrt{nN}}\frac{1}{n}\sum_{j}\overline{\mathfrak{b}}_{ij}[\widehat{u}_{\eta}]\,\overline{y}_{ij}^{2}.$ (A.74)
Hence,  [^1] Recalling (A.30), we also have for any $\mathsf{c}>0$, $\max_{i\leq N}|h^{\mathsf{c}-1/2}(\widehat{\tau}_{i}-\tau_{i})|\lesssim(1+N^{-1/2}|\widehat{\eta}|)\max_{i\leq N}|h^{\mathsf{c}}\,\overline{\mathfrak{m}}_{i}|=O_{p}(1)o_{p}(1)=o_{p}(1)$.
$\left|\sqrt{n}(\widehat{\tau}_{i}-\tau_{i})\right|\lesssim|\overline{\mathfrak{m}}_{i}|+\frac{|\widehat{u}_{\eta}|}{\sqrt{N}}\frac{1}{n}\sum_{j}\overline{y}_{ij}^{2}\lesssim|\overline{\mathfrak{m}}_{i}|\left(1+\frac{|\widehat{u}_{\eta}|}{\sqrt{N}}\right).$ (A.75)
@@
Since $\widehat{u}_{\eta}=O_{p}(1)$, we can take $K>0$ large enough to ensure that $\sup_{n,N}P[A_{12,1}^{c}]<\epsilon^{\prime}$. With this $K$, by (A.75) and the Markov inequality we have for any $L\geq 2$,
@@
Finally, by Assumption 2.4, it is easy to see that
$P\left[A_{12,1}\cap A_{12,2}\cap\{\delta_{12}>\epsilon\}\right]\leq P\left[\sqrt{\frac{N}{n}}\,\frac{1}{N}\sum_{i=1}^{N}|\overline{\mathfrak{m}}_{i}|>\epsilon\right]\lesssim\sqrt{\frac{N}{n}}\to 0.$ (A.79)
@@ A.2.3. Joint asymptotic normality
Having obtained the asymptotic linear representations (A.64) and (A.71), it remains to show the asymptotic orthogonality, hence the asymptotic independence of $\widehat{\eta}$ and $\widehat{\theta}_{\tau}$. We only need to observe that the covariance tends to 0.
By what we have seen in Sections A.2.1 and A.2.2, in particular the stochastic expansions (A.65) and (A.71), the computation of the covariance of the $\Delta_{11}$ and $\Delta_{12}$ boils down to that of
${\rm Cov}[\zeta_{i},\zeta_{i}^{\prime}]$ $=\frac{1}{2}\,{\rm Cov}\left[M_{11,1}-\overline{g}_{1}M_{11,0},\,\partial_{\theta_{\tau}}\log f(\tau_{1};\theta_{\tau,0})\right].$ (A.80)
Indeed, it can be seen that $\lim_{n,N}{\rm Cov}[\zeta_{i},\zeta_{i}^{\prime}]=0$ ($p_{\eta}\times p_{\tau}$-zero matrix) in a similar manner to the previous $\mathcal{F}$-stable convergence argument of $(M_{11,0},M_{11,1})$ combined with the uniform integrability that we have used to obtain the limit of the expectation in (A.54). This completes the proof of Theorem 3.2.
@@
To study the asymptotic behavior of second-stage marginal quasi-likelihood $\mathbb{H}_{2}(\vartheta_{2})$ defined by (3.17), we will proceed as follows.
  1. (i)
@@
    for specific leading terms $M_{i,T}$ and $v_{i,T}$ (recall their definitions in (3.19)) and some remainder (negligible) terms $\mathsf{R}^{M}_{i}$ and $\mathsf{R}^{v}_{i}$.
  2. (ii)
@@
    $\mathbb{H}^{\star}_{2}(\vartheta_{2})$ $:=\sum_{i}\log\phi_{p_{\varphi}}\big(M_{i,T}^{-1}v_{i,T};\,\mu,\,M_{i,T}^{-1}+\Sigma\big)$
    $=\sum_{i}\log\phi_{p_{\varphi}}\big(M_{i,T}^{-1}v_{i,T};\,(\varphi_{f},0),\,M_{i,T}^{-1}+\mathrm{diag}(O,\Sigma_{r})\big)$ (A.82)